OpenStudy (anonymous):

Given that x and y are positive integers, solve the equation x^2-4y^2=37

Directrix (directrix):

Mr. Wolf cranked out ± 19 for x and ± 9 for y. Look at the graph - it is a hyperbola. Follow the link: http://www.wolframalpha.com/input/?i=x%5E2+-+4y%5E2+%3D+37

OpenStudy (anonymous):

But how do I show the working?

Directrix (directrix):

I'll begin the solution for x and you can afterwards do so for y. Before that, though, the question states that x and y are positive integers, so the answers are 19 for x and 9 for y. Those are answers and not solutions. The solution is comprised of the answer and how it was obtained.

OpenStudy (nikvist):

\[x^2-4y^2=(x-2y)(x+2y)=37=1\cdot 37\]\[x-2y=1\quad,\quad x+2y=37\quad\Rightarrow\quad x=19\quad,\quad y=9\]

OpenStudy (anonymous):

But how do I know that x-2y is equal to 1 while x+2y is equal to 37? :O

OpenStudy (nikvist):

You have two choices: 1.\[x-2y=1\quad,\quad x+2y=37\quad\Rightarrow\quad x=19\quad,\quad y=9\]2.\[x-2y=37\quad,\quad x+2y=1\quad\Rightarrow\quad x=19\quad,\quad y=-9\not\in\mathbb{N}\]

OpenStudy (anonymous):

Ok. Thank you so much!